OpenStudy (anonymous):

Here is my problem: (7m^9)^3 So, I believe I would do this to continue the problem. = 7^3(m^9)^3 = 7^3 m^27 = 7m^27 Is this correct? If not, please explain.

OpenStudy (anonymous):

ur last step is wrong.

OpenStudy (anonymous):

in the second step u got 7^3 which is correct but then in your last step u have only 7.

OpenStudy (anonymous):

do u know where u went wrong now?

OpenStudy (anonymous):

tell me ur answer and I'll check it for u.

OpenStudy (anonymous):

Oh, I think it is because the ^3 cancels out on the 7 right?

OpenStudy (anonymous):

nope, how can it cancel out on the right?

OpenStudy (anonymous):

That's true... can you explain it?

OpenStudy (anonymous):

k wait, I am having problems with the post.

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

remember to give me a medal lol :P

OpenStudy (anonymous):

Okay haha

OpenStudy (anonymous):

\[(7m^9)^{3}\] \[=7^{3}m^9(^{3})\] =7^3m^27

OpenStudy (anonymous):

\[(7^1m^9)^{3}\] \[7^3m ^{27}\] That's all.

OpenStudy (anonymous):

whenever there are powers/exponents outside ( )^3 just times the exponent by the exponent of the terms inside.

OpenStudy (anonymous):

Oh, alright. I get it now. Thank you!

OpenStudy (anonymous):

Explanation: There is nothing to cancel you just times the powers.

OpenStudy (anonymous):

no problem and u didn't forget my reward lol :P

OpenStudy (anonymous):

Nope, I didn't haha